OpenStudy (anonymous):

integral of 7/ (e^t)^7 Okey this is how I broke it down:

OpenStudy (anonymous):

\[u = e^{7t} wich gives du = 7e^{7t}\] \[\frac{ du }{ dt } = 7e^{7t} which gives \frac{ du }{e ^{7t} } = 7 dt\] and then \[\int\limits_{}^{} \frac{ 1 }{ e^{7t} } * 7\] can be replaced with \[\int\limits_{}^{} \frac{ 1 }{ u } * \frac{ du }{ e^{7t} }\]

OpenStudy (anonymous):

So far is it right? if not the rest is probably wrong and I should wait for this to clear up.

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ 7 }{ (e^{t})^7 }\] is the integral to make it clear.

OpenStudy (anonymous):

7/(e^t)^7=7/e^7t =7e^(-7t)

OpenStudy (anonymous):

So you can just integrate directly

OpenStudy (anonymous):

you can see my file....

OpenStudy (anonymous):

if you integrate 7e^(-7t) you will get -e^(-7t)+C

OpenStudy (anonymous):

right...

OpenStudy (anonymous):

-e^-7t + c is the correct answer. song6 I couldn't quite follow your equation.

OpenStudy (anonymous):

alright with dominics way it was very easy since I could avoid all the steps, alright, thanx to both for the help and time:) closing the thread now.